### v1

# UnZipLoRA: Separating Content and Style from a Single Image

Abstract
* Currently at Google.   Currently at Meta.
   Project page: https://unziplora.github.io
222
This paper introduces UnZipLoRA, a method for decomposing an image into its constituent subject and style, represented as two distinct LoRAs (Low-Rank Adaptations). Unlike existing personalization techniques that focus on either subject or style in isolation, or require separate training sets for each, UnZipLoRA disentangles these elements from a single image by training both the LoRAs simultaneously. UnZipLoRA ensures that the resulting LoRAs are compatible, i.e., they can be seamlessly combined using direct addition. UnZipLoRA enables independent manipulation and recontextualization of subject and style, including generating variations of each, applying the extracted style to new subjects, and recombining them to reconstruct the original image or create novel variations. To address the challenge of subject and style entanglement, UnZipLoRA employs a novel prompt separation technique, as well as column and block separation strategies to accurately preserve the characteristics of subject and style, and ensure compatibility between the learned LoRAs. Evaluation with human studies and quantitative metrics demonstrates UnZipLoRA’s effectiveness compared to other state-of-the-art methods, including DreamBooth-LoRA, Inspiration Tree, and B-LoRA.
Figure 1: Given a single input image (middle), UnZipLoRA learns a disentangled content or subject LoRA (left) and style LoRA (right) that can be used to generate new images with the learned concepts.
1 Introduction
Imagine an artist inspired by a single image – perhaps a microscope rendered in a unique illustration technique (Fig. 1). This image depicts both the content or subject (the microscope with its specific shape and details) and style (the characteristic artistic technique). What if we could disentangle these intertwined elements, enabling the artist to extract and manipulate them independently? This capability, long sought in the computer vision literature [36], would open up new avenues for artistic expression, allowing for the destylization and recontextualization of the subject in different conditions, the application of the extracted style to new subjects, and the creation of entirely novel combinations while preserving both the original subject and style.
The recent surge of diffusion models [13, 35, 27] has unlocked previously unprecedented automatic image generation capabilities. The primary means of controlling such models is through text prompts, but text-based conditioning is inadequate for capturing the details of nuanced concepts, such as specific object instances (my dog) or individual styles (my child’s crayon drawing). Example-driven generation is highly desired in such scenarios. To this end, model fine-tuning methods like DreamBooth [29] and StyleDrop [33] captures subject or style of the reference image(s), and allow for the generation of novel renditions of specific subjects or styles. However, such approaches tend to focus on either content or style in isolation, and cannot be easily made to capture both or perform disentanglement. On the other hand, stylization techniques like ZipLoRA [32] and B-LoRA [6] can combine the subject and style together, but require the training of two separate models using subject- and style-specific input images. However, scenario we consider requires the opposite,i.e., decomposing the subject and style from a single input image.
This challenge of separating content and style from a single image is precisely what we tackle in this work. We introduce UnZipLoRA, a novel method that deconstructs an image into its constituent subject and style, represented as two distinct LoRAs (Low-Rank Adaptations [14]) trained simultaneously. These LoRAs can be used independently to generate variations of the subject or style, and allow for recontextualizations. Moreover, our joint training method ensures that the resulting LoRAs are inherently compatible,i.e., can be seamlessly combined by direct addition to reconstruct the original image or to generate novel compositions of subject and style while preserving their fidelity. In fact our approach can be seen as a next stage of concept extraction - a problem previously studied in concept decomposition methods like Inspiration Tree [37] and CusConcept [40] that rely on textual inversion [8] to learn multiple text embeddings corresponding to the subconcepts within a set of images. However, textual inversion alone, without weight fine-tuning, does not provide adequate control over or fidelity of the extracted concepts that remain generic in nature and do not capture nuances of the input object/style.
As suggested by the name, UnZipLoRA operates in the opposite direction of ZipLoRA [32], which is focused on merging independently trained subject and style LoRAs. While ZipLoRA addresses the challenge of combining pre-existing LoRAs, UnZipLoRA tackles the inverse problem: disentangling a single image into its subject and style components such that the resulting LoRAs can be used either together or separately. Mathematically speaking, the decomposition problem is ill-posed and cannot be trivially derived from the approach of ZipLoRA [32].
Our key challenge is to learn two independent LoRAs simultaneously using only a single input image as supervision while ensuring that the resulting LoRAs correctly capture the subject and style concepts. Typical LoRA fine-tuning operates by binding the LoRA weights with the trigger phrase representing a specific subject or style concept in the input image – such as “a <$c$> dog in <$s$> painting style”. If we apply such a naive approach in our case, the presence of both the trigger phrases in a single input prompt makes it difficult for the subject and style LoRAs to bind to the correct trigger phrase, resulting in cross-contamination. To solve this problem, we propose a novel prompt separation strategy that uses different prompts for each LoRA and the base model, and then combines them together in the intermediate feature space of the diffusion model in such a way that the loss for each LoRA can be calculated jointly using only the input image as supervision.
While prompt separation allows for joint training of LoRAs, the resulting LoRAs may not be compatible with each other, i.e., combining them through direct addition may produce poor quality recontextualizations. To make them inherently compatible, we also propose column separation and block separation strategies built upon insights from ZipLoRA [32] and B-LoRA [6] respectively. Column separation determines the importance of each column of LoRA weight matrix and dynamically assigns each column to either a subject LoRA or style LoRA. Such disjoint assignment ensures that high-importance columns from each LoRA are not added together. Block separation reserves some blocks of the U-net predominantly for style or for the subject, ensuring a further degree of disentanglement.
We demonstrate the effectiveness of UnZipLoRA for accurate separation of content and style using both human studies and quantitative metrics. Our results show a clear advantage of our method over separate LoRA fine-tuning via DreamBooth [29], concept separation via Inspiration Tree [37], and even the most recent state-of-the-art B-LoRA method [6]. We also showcase our method’s ability to preserve the concept fidelity for wide array of recontextualizations – whether for using subject or style separately or together.
2 Related Work
Fine-tuning diffusion models is an effective way to personalize the text-to-image (T2I) models to depict specific concepts based on textual descriptions. Textual Inversion [8] optimizes the text embedding to represent a specific visual concept. DreamBooth [29] fine-tunes the diffusion model itself to better capture an input concept from a small number of images. Similarly, another group of methods [10, 17] aims to optimize specific parts of the networks to capture visual concepts. Importantly, LoRA [14] is a finetuning techique that only optimizes a small subset of weights by low-rank approximations. Personalization approaches quickly adopt this techique, as it is efficient for training and can mitigate overfitting problems.
Image stylization is an area of research dating back at least 20 years [5, 12]. Great advances in arbitrary style transfer was achieved by the convolutional neural network-based approaches [9, 16, 15, 19, 24]. With the advent of image generative models, a variety of approaches have attempted to fine-tune a pre-trained generative model such as GAN or diffusion model for the task of stylization [20, 7, 43, 23, 38, 18, 3, 31, 42, 41, 33, 4]. While these works provide valuable insights into style learning using generative models, the task we attempt to solve in this work is the opposite: instead of stylizing a content image, we attempt to decompose a stylistic image into its subject and style counterparts.
Content-style decomposition. The task of subject-style decomposition can also be seen as a type of concept extraction. Inspiration Tree [37] aims to learn multiple embeddings corresponding to hierarchical subconcepts within a set of images. However, its reliance on textual inversion [8] limits this method to primarily producing text embeddings rather than the full LoRA weights needed for granular control over generation. Lego [22] is another method based on textual inversion that explores inverting personalized concepts from images that go beyond object appearance and style. Similarly, CusConcept [40] decomposes an image into visual concepts by learning customized embeddings. However, it does not produce dedicated LoRA models for each concept. This limitation, along with its reliance on computationally expensive LLMs and its lack of explicit content-style decomposition, restricts its practical applicability and flexibility. U-VAP [39] is a fine-tuning method that allows users to specify desired attributes from a set of images, enabling the disentangled use of visual concepts in diverse settings. However, this method requires elaborate data augmentation aided by LLMs, posing scalability challenges. ConceptExpress [11] explores unsupervised concept extraction and recreation by leveraging the inherent capabilities of pre-trained diffusion models. However, its reliance on localized masks extracted from the U-net limits its application to concepts that can be localized within an image. This constraint prevents it from effectively handling global concepts like style. This limitation applies to Break-A-Scene [1] too, which relies on localized masks.
The method most directly aimed at content and style separation is B-LoRA [6]. The authors of this work analyze the architecture of the SDXL base model [25] to find blocks most responsible for capturing content and style. By jointly training content and style LoRAs restricted to the respective blocks, they obtain models that can be successfully mixed and matched for various stylization applications. We take advantage of the same block constraints in our work, but find that they are not sufficient. Tellingly, B-LoRA [6] does not demonstrate results for de-stylization – i.e., given a stylized image, render the same subject in a realistic style. In our experiments, we find that B-LoRA struggles to extract and represent subjects in their unstylized appearance, hindering its ability to faithfully manipulate content. The prompt separation strategy of UnZipLoRA provides a significant improvement in disentanglement ability, and column separation further enhances the fidelity of fine details.
3 Method
Figure 2: Overview. (1): Our training pipeline takes three prompts to learn two separate LoRAs. The column masks are introduced to establish the orthogonality. (2)-(4): Inference with individual learned LoRA or the combined LoRA.
3.1 Preliminaries
Diffusion models.  [13, 35, 27] represent the current state-of-the-art in generative modeling, renowned for their ability to synthesize high-quality, photorealistic images. The forward diffusion process gradually adds Gaussian noise to an image until it becomes indistinguishable from random noise, while the training stage learns a model for the reverse process that reconstructs the original image by gradually removing noise. This reconstruction is typically achieved using a U-net architecture with text conditioning, enabling text-to-image generation. In this work, we focus on latent diffusion models (LDMs) [27], which operate in a lower-dimensional latent space. Specifically, we utilize Stable Diffusion XL v1 [25] for all our experiments.
LoRA fine-tuning. LoRA (Low-Rank Adaptation) is a method for efficient adaptation of large vision or language models to a new downstream task [14, 30]. The key concept of LoRA is that the weight updates $\Delta W$ to the base model weights $W_{0}\in\mathbb{R}^{m\times n}$ during fine-tuning have a low intrinsic rank, thus the update $\Delta W$ can be decomposed into two low-rank matrices $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ for efficient parameterization with $\Delta W=BA$. Here, $r$ represents the intrinsic rank of $\Delta W$ with $r<<min(m,n)$. During training, only $A$ and $B$ are updated to find suitable $\Delta W=BA$, while keeping $W_{0}$ constant. For inference, the updated weight matrix $W$ can be obtained as $W=W_{0}+BA$. Due to its efficiency, LoRA is widely used for fine-tuning open-sourced diffusion models. In case of text-to-image LDM models, model customization can be achieved by fine-tuning the LoRA to minimize reconstruction loss $\mathcal{L}_{DB}$ as proposed by DreamBooth [29].
3.2 Problem Setup
We aim to decouple content and style from a single input image by learning two distinct LoRAs simultaneously. Given a pre-trained diffusion model with weights $\{W_{0}^{i}\}$, we learn two models: content LoRA $L_{c}=\{\Delta W_{c}^{i}\}$ and style LoRA $L_{s}=\{\Delta W_{s}^{i}\}$. Here, $i$ denotes the index of layers of the diffusion U-net. We use the trigger phrases <$c$> and <$s$> in our text prompts to describe the content and style respectively. Once trained, the resulting LoRAs can be used either separately or together to achieve various recontextualizations as depicted in Fig. 2. For simplicity, we drop the superscript $i$ since our method operates over all the LoRA-enabled weight matrices of the base model.
3.3 UnZipLoRA
In this section, we propose our UnZipLoRA approach, which relies on three key components to ensure accurate disentanglement. First, prompt separation is introduced in Sec. 3.3.1 to train the two LoRAs with different prompts emphasizing style or content. Second, column separation is introduced in Sec. 3.3.2 to enforce the orthogonality of weights in the two LoRAs with the help of a sparse column map. Third, we introduce block separation in Sec. 3.3.3 to assign more weight to the style LoRA and content LoRA in the style-sensitive or content-sensitive blocks, respectively.
3.3.1 Prompt Separation
Our core challenge lies in the utilization of a single image as supervision for training two separate LoRAs. Since the input image contains both subject and style, we must use both LoRAs $L_{c}$ and $L_{s}$ together to calculate the loss during training. If we also use <$c$> and <$s$> in the input prompt, as in “A <$c$> in <$s$> style”, the training would lead to cross-contamination since the tokens corresponding to the style descriptor <$s$> would be attended to by the cross-attention layers of the subject LoRA, and vice versa. This is evident in Fig. 3, where both the subject and style LoRAs are overfitted to input image when a naive DreamBooth-LoRA is used.
Cross-attention layers in diffusion models are responsible for learning the text conditioning, and play a crucial role in binding the target concepts to the corresponding parts of the prompt. In a typical cross-attention layer of the diffusion U-Net, the prompt embedding $x$ is mapped to keys $K$ and values $V$ in the transformer using weights $W_{0}$:
$K(x)~{}\textnormal{or}~{}V(x)=W_{0}^{T}x\,.$ (1)
If we add the content and style LoRAs to the base model and use both <$c$> and <$s$> in the prompt $x$, the mapping in the cross-attention layer becomes
$K(x)~{}\textnormal{or}~{}V(x)=\left(W_{0}+\Delta W_{c}+\Delta W_{s}\right)^{T}x\,.$ (2)
This allows the content (resp. style) LoRA to attend to style (resp. content) tokens, resulting in cross-contamination. Instead of the naive strategy in eq. (2), we propose to calculate three sets of keys and values using three separate prompts: one with the base model $W_{0}$ and combined prompt $x$, and one each with the content and style LoRA and their respective prompts (See Fig. 2). The resulting feature maps are then added together as
$K~{}\textnormal{or}~{}V(x,x_{s},x_{c})=W_{0}^{T}x+\Delta W_{s}^{T}x_{s}+\Delta W_{c}^{T}x_{c},$ (3)
where $x$ is the embedding for the combined prompt “A <$c$> in <$s$> style”, while $x_{c}$ and $x_{s}$ are the embeddings of subject and style descriptors <$c$> and <$s$> respectively. This allows each LoRA to attend to different concepts, preventing cross-contamination as shown in Fig. 3: prompt separation successfully destylizes the content along with bringing improvements in style learning.
Figure 3: Effectiveness of each component of our approach. We use DreamBooth to train two LoRAs with Eq. 2 as our baseline. The prompt separation achieves successful decomposition of the concepts into different LoRAs. The column separation disentangles the two LoRAs to make them compatible for joint inference. The block separation further boosts the quality of the learned style.
3.3.2 Column Separation
Apart from learning separate concepts, we want our resulting LoRAs to be compatible, i.e. we want to be able to combine them through direct arithmetic merge to generate the subject and style together. Prompt separation effectively guides the two LoRAs to learn distinct concepts, and especially helps in achieving better destylization of the content. However, it does not guarantee compatibility since training LoRAs with different prompts can result in weight misalignment when they are combined to process the same prompt during inference. As seen in Fig. 3, output of the combined subject+style LoRA does not adhere the input prompt and overfits to the input image.
To address this, we introduce the concept of column masks for each LoRA, denoted as $m_{s}$ and $m_{c}$. These column masks dynamically control the contribution of each column in the learned LoRA weights (see Fig. 2). Essentially, they allow the model to selectively activate or suppress specific columns within each LoRA, promoting orthogonality and reducing interference. By incorporating these column masks, the attention block update is modified as follows:
$K~{}\textnormal{or}~{}V(x,x_{s},x_{c})=W_{0}^{T}x+m_{s}\Delta W_{s}^{T}x_{s}+m_{c}\Delta W_{c}^{T}x_{c}.$ (4)
Orthogonal loss. To promote compatibility between the subject and style LoRAs, we leverage the insight from ZipLoRA [32] that highly aligned LoRA columns merge poorly, and apply the following loss to $m_{c}$ and $m_{s}$:
$\vspace{-0.2em}\mathcal{L}_{\perp}=\sum_{i}|m_{c}^{i}\cdot m_{s}^{i}|,\vspace{-0.2em}$ (5)
minimizing which promotes orthogonality between the learned content and style weights. Orthogonal loss $\mathcal{L}_{\perp}$ is added to our reconstruction loss $\mathcal{L}_{DB}$ as regularizer with the weight parameter $\lambda_{\perp}$.
Figure 4: Qualitative Comparisons. We compare subject and style decomposition outputs from UnZipLoRA against DreamBooth-LoRA [29], Inspiration Tree [37], and B-LoRA [6]. Our results demonstrate that UnZipLoRA achieves superior subject and style fidelity, effectively disentangling and preserving these concepts compared to alternative approaches.
Sparse masks with importance re-calibration. To further decompose the concepts, instead of training the entire LoRA matrices, we find that training 30% of the columns in each weight matrix is sufficient to learn concepts, and it ensures weight sparsity for a good decomposition. This strategy is inspired by Liu et al. [21], who find that a small set of neurons tend to be much more salient than the others for capturing concepts.
We use a dynamic approach to select the most important 30% of the columns during training. Before the training, we initialize the column masks, $m_{s}$ and $m_{c}$, with the top 10% of the most important columns for style and content. Importance is calculated using the Cone method [21], with five warm-up training steps using the full LoRA weights. Then, during the remainder of training, for every $t$ steps, we re-calibrate the column masks by calculating the column importance of LoRA weights and adding the new top 10% of the most important columns to $m_{s}$ and $m_{c}$ until the 30% cap is reached. Our strategy brings significant improvements in compatibility of the resulting LoRAs, preventing overfitting and producing better recontextualization as shown in Fig. 3.
3.3.3 Block Separation
Furthermore, we take advantage of the analysis by the authors of B-LoRA [6], who find that certain blocks in the SDXL U-Net are more responsible for content or style. To leverage this prior knowledge, for blocks known as style-sensitive, we release the column sparsity constraint for the style LoRA,i.e., all the columns are fully trained. The same strategy is applied to the content-LoRA. This further improves the accuracy of fine details, especially for the style LoRA (see Fig. 3).
4 Experiments
4.1 Implementation Details
Dataset. For our experiments, we use a diverse set of $40$ stylistic images representing unique styles and subjects as input images. These images are sourced from previous works, online repositories offering free-to-use content, and from state-of-the-art text-to-image generators. The attribution and license information of the input images used in our experiments is provided in the supplementary.
Experimental setup. We use SDXL v1.0 [25] as our base model for all our experiments. We use rank$=64$ for our subject and style LoRAs, and fine-tune them using Adam optimizer with learning rate of $55e-$ 5. Our LoRAs are trained for $600$ steps using batch size of $1$. We keep the base SDXL model and the text encoders frozen during training. For column separation, we choose $t=200$ as the sparse column re-calibration frequency. For orthogonal loss regularization, we use $\lambda_{\perp}=0.5$ in all our experiments.
While B-LoRA inspires our block separation strategy, we diverge significantly in selecting blocks for subject and style representation. Instead of utilizing only two distinct blocks from the upsampling blocks of the U-Net, as in B-LoRA, our method leverages all upsampling blocks within the SDXL architecture for a comprehensive representation of both subject and style. We partition these upsampling blocks and assign a subset of them for content learning while the remainder is used for style learning. A detailed description of our block partitioning strategy is provided in the supplementary.
Figure 5: Recontextualization. The trained style LoRA and content LoRA can be used individually or jointly at the inference time. The learned concepts can be used to generate images in various contexts, validating our method’s robustness.
Our choices of training prompts are standard and inspired from previous works [29, 33, 32]. Our combined prompt is of the form “a <$c$> in <$s$> style”, while “A <$c$> ” and “An image in <$s$> style” are used as subject- and style-specific prompts respectively where <$c$> and <$s$> are the subject and style descriptors. As per previous works [29, 33, 32], we use subject category along with unique token (such as “sks”) to describe the subject, and use detailed artistic description for the style. For example, the prompt we use for the first input image in Fig. 4 is “A sks piano in abstract impressionism style”. More such implementation details are deferred to supplementary.
4.2 Subject and Style Decomposition
We take a single input image and apply our UnZipLoRA algorithm to obtain the style and subject LoRAs. We use these LoRAs separately to generate subject-only and style-only representations of the input image with diverse text prompts. The subject and style separation results obtained by our approach are shown in Fig. 1 and Fig. 4 from various input images. The quality of the learned concepts is further validated with various recontextualization examples in Fig 5. Our method successfully captures the intricate details of both the style and subject and can effectively disentangle both elements, as validated by the quantitative and qualitative evaluations.
Qualitative comparisons. We present qualitative comparisons of our approach with other methods in Fig. 4. We use DreamBooth-LoRA [29] as our baseline – we train subject and style LoRA models separately using the DreamBooth method with the “A <$c$> ” and “A subject in <$s$> style.” While DreamBooth-LoRA demonstrates satisfactory performance in learning the style, it fails to effectively destylize the subject and results in overfitting the input image. This highlights the inherent difficulty in disentangling subject and style from a single image, even when training LoRAs independently.
UnZipLoRA is next compared with Inspiration Tree [37], a textual inversion-based concept extraction approach. Inspiration Tree aims to separate the properties of the input into a binary-tree-structured concept tree using Textual Inversion [8]. We train their model to separate subject and style with the prompt “A <$c$> in <$s$> style”, initializing each placeholder with “object” and “art”. Despite employing separate training for content and style, Inspiration Tree struggles to distinguish and disentangle these properties accurately. While it could identify the overarching category of concepts correctly, it fails to capture the intricate details of the subject or the style. This limitation stems from its reliance on Textual Inversion, which focuses on learning text embeddings rather than fine-tuning the weights, resulting in limited expressiveness and controllability.
Figure 6: Comparison of Subject-Style Recontextualization. We present a comparison of subject-style recontextualization between our method and B-LoRA across diverse prompts and input images. The results highlight our method’s superior ability to flexibly adapt subjects and styles to various contexts while accurately reproducing both subject and style features.
We also compare our approach with B-LoRA [6] – a technique aimed at combining style and the subject of two different images. To adapt B-LoRA to our problem setting, we utilize the same input image to train both the subject and style representations while keeping other hyperparameters, such as the learning rate and training steps, consistent with those reported in their paper. UnZipLoRA outperforms B-LoRA in several key aspects. Firstly, UnZipLoRA demonstrates more stable content destylization and better consistency in generating variations. This is attributed to our prompt separation strategy, which ensures that each LoRA focuses exclusively on its designated concept, promoting accurate disentanglement. Secondly, UnZipLoRA exhibits superior preservation of content fidelity.
As evident in the mushroom and balloon examples in Fig. 4, B-LoRA results often include residual style features and extraneous background elements, indicating overfitting and incomplete disentanglement. Furthermore, B-LoRA fails to accurately capture the color of the balloon, treating it as part of the style rather than the content. In contrast, UnZipLoRA effectively mitigates overfitting through our column separation strategy, which employs orthogonal column masking to ensure compatibility of resulting LoRA weights. The limitations of B-LoRA can be attributed to its block-wise training approach, where content information is confined to specific blocks of the U-Net, potentially leading to information loss. Moreover, unlike B-LoRA, our method trains both the LoRAs jointly, reducing the compute resource requirements significantly. More quantitative evaluation results can be found in the supplementary.
Quantitative results. For quantitative comparisons, We conduct user studies comparing our method with the competing approaches. In our study, each participant is shown the input image, along with the outputs of the two methods being compared. Each output group consists of $4$ images for the subject and $4$ images for style selected randomly. Participants are asked to choose an output group that decomposes the input image into style and content more accurately. We conducted three separate user studies: one for each competing method vs. UnZipLoRA. We received $204$ responses for each study from $34$ total participants. As can be seen in Tab. 1 UnZipLoRA receives higher user preferences than all three competing methods owing to its superior disentanglement capabilities.
Table 1: User Preference Study. We compare the user preference of accurate stylization and subject fidelity between our approach and competing methods. Users generally prefer our approach.
[…]
DreamBooth LoRA Inspiration Tree B-LoRA
$91.17\%$ $81.53\%$ $62.74\%$
Table 2: Subject-alignment and Style-alignment Scores. We measure the cosine similarities between the input image and the decomposed output using CLIP-I and DINO features for style and subject respectively. We also provide alignment scores using CSD [34], a model more suitable for stylistic images. As indicated by higher alignment scores, UnZipLoRA provides superior subject and style fidelity on all the metrics.
DreamBooth -LoRA Inspiration Tree B-LoRA UnZipLoRA
Style-align. (CLIP-I) $\uparrow$ $0.417$ $0.404$ $0.418$ $\mathbf{0.427}$
Subject-align. (DINO) $\uparrow$ $0.339$ $0.291$ $0.337$ $\mathbf{0.349}$
Style-align. (CSD) $\uparrow$ $0.245$ $0.229$ $0.244$ $\mathbf{0.265}$
Subject-align. (CSD) $\uparrow$ $0.338$ $0.334$ $0.342$ $\mathbf{0.358}$
Figure 7: Ablation Examples. We show the impact of adding prompt-wise separation, column-wise separation, and block-wise separation sequentially. Each row illustrates the outputs of baseline methods and our proposed approaches, highlighting their difference in subject-style disentanglement, style fidelity, and recontextualization across different examples.
We also provide quantitative comparisons of our approach using subject-alignment and style-alignment scores in Tab. 2. To calculate the alignment scores, we start with commonly used CLIP-I image embeddings for style and DINO features for objects. We calculate feature embeddings for our style and subject decomposition using CLIP-I [26] and DINO [2] features, respectively, using the input image as the reference. In both cases, we use cosine similarity as the metric and calculate averages over $8$ input images by generating $16$ samples for each. As shown in Tab. 2, UnZipLoRA achieves the highest alignment scores for both the subject and style, highlighting its superiority. DreamBooth-LoRA tends to overfit the input image, resulting in higher scores even though its performance is inferior to other methods, as depicted in qualitative results.
One should note that these metrics are inherently limited, particularly in measuring style alignment, since they may not fully capture subtle stylistic nuances and can be influenced by the semantic content of the images. To this end, we use a recently proposed CSD model [34] trained specifically to extract the style descriptors from images. We use the embeddings from CSD’s content and style branches to measure subject and style fidelity respectively. The CSD alignment scores (Tab. 2) indicate a clearer separation between our method and the others, highlighting the superiority of UnZipLoRA.
Subject-style combination and recontextualization. A key advantage of UnZipLoRA is its ability to produce compatible subject and style LoRAs that can be seamlessly merged via direct addition. This allows for the generation of novel recontextualizations that faithfully incorporate both the subject and style of the original image. We demonstrate this capability in Fig. 5 through various recontextualizations, showcasing the versatility of UnZipLoRA in generating images using individual LoRAs or a combination of both. This capability to manipulate and recombine the extracted concepts underscores the robustness of our method and its potential for creative applications.
We also present additional examples of subject-style recontextualization compared with B-LoRA, which is specifically designed for it, using diverse prompts and input images in Fig. 6. The results demonstrates that our method effectively enables flexible and accurate adaptation of subjects and styles to different contexts, outperforming B-LoRA. Our generations match the requested contexts while preserving both subject and style features.
As noted in subject-only and style-only comparisons, B-LoRA suffers from overfitting and inconsistent feature preservation, issues that also affect subject-style generation. For instance, in the sun hat example of Fig. 6, B-LoRA fails to retain key features of the input hat, demonstrating these limitations. Moreover, our method demonstrates clear superiority over B-LoRA in handling diverse scenarios, including those involving multiple subjects or complex relationships. For instance, in the pizza example in Fig. 6, B-LoRA incorrectly applies subject features to the cat, generating pizza-like textures. In contrast, our method perfectly preserves the cat and the pizza, generating a contextually appropriate output.
4.3 Ablation Examples
In this section, we present qualitative examples to illustrate the impact of the different components of our method. We use DreamBooth [29] as the baseline, and iteratively add prompt-, column-, and block-separation strategies. We compare the performance of each ablated version of our method to demonstrate effectiveness of each component. For simplicity, let us refer to ablations of our method as M1, M2, and M3, where:
  • M1 is Baseline + Prompt Separation
  • M2 is Baseline + Prompt + Column Separation
  • M3 is our full UnZipLoRA method (Baseline + Prompt + Column + Block Separation)
We provide a qualitative comparison of the effects of each proposed component in Sec. 3.3 in Fig. 7. Each row demonstrates the recontextualization of subject, style and combined subject-style across different configurations.
As shown in Fig. 7, adding only the prompt-wise separation (M1) to the baseline significantly improves the disentanglement of content with style compared to baseline. However, M1 still struggles to capture the style complexities from the input images, and shows strong overfitting to the input image in subject-style recontextualization highlighting the incompatibility of resulting LoRAs.
Next, as described in detail in Sec. 3, incorporating column-wise separation (M2) reduces the interference between subject and style LoRAs and improves their compatibility while preserving subject and style generation abilities. For example, details like the color of the microscope in Fig. 7 are retained, while the combined recontextualization performance improves significantly. However, improvements in style decomposition remain modest, indicating that using a small portion of the columns for style is insufficient for effective style learning.
Furthermore, we introduce block separation (M3) with subject- and style-specific blocks, further refining the disentanglement of subject and style. This addition significantly enhances the model’s ability to capture fine-grained stylistic features owing to dedicated style-learning blocks. Fig. 7 shows that all three separation strategies together achieves the best results across all examples, enabling comprehensive disentanglement of subject and style and flexible recontextualization.
4.4 Limitations
In some cases involving abstract artistic styles, UnZipLoRA may not achieve perfect destylization. For instance, in Fig. 8, the kangaroo example exhibits artifacts on its leg due to the influence of the curved line-art style, while the bear’s form appears slightly unnatural due to the high level of abstraction in the original style. This highlights the challenge of disentangling content from highly abstract styles where the style significantly influences the subject’s depiction.
Figure 8: Failure Cases. For a few input images with abstract styles, UnZipLoRA fails to accurately destylize the subjects. Note the artifact introduced on kangaroo’s leg due to curved line-art style, and unnatural shape of the bear due to high levels of abstraction.
5 Conclusion and Future Work
We introduced UnZipLoRA, a novel method for decomposing a single image into its subject and style, represented as two disentangled and compatible LoRAs. By employing prompt, column, and block separation strategies, UnZipLoRA effectively extracts and preserves these elements, enabling diverse recontextualizations and manipulations. Our experiments demonstrate superior performance compared to existing methods, highlighting UnZipLoRA’s potential for creative exploration and control within text-to-image generation. Future work includes exploring alternative disentanglement techniques and training-free approaches for subject-style decomposition.
Supplementary Material
A Additional Implementation Details
In this section, we provide additional implementation details for our algorithm:
Block separation strategy. As discussed in Sec. 3.3.3, we employ block separation technique similar to that proposed in B-LoRA. Specifically, as shown in Fig. 9, the U-Net architecture in SDXL comprises three primary components: the downsampling blocks, the middle blocks, and the upsampling blocks. Each of these components contains several groups of transformer-based blocks. As the upsampling component is more critical to the overall performance, our primary focus for block separation lies in upsampling. Within the upsampling component, there are two distinct groups of blocks, which are differentiated by number of transformers per block: the first group (Upblock0 in Fig. 9) contains blocks with 10 transformers each, while the second group (Upblock1) has blocks with only 2 transformers. Due to this disparity, the first group plays a more significant role in the learning process.
Figure 9: Block separation strategy in the diffusion U-Net architecture of SDXL. The illustration highlights the separation of transformer blocks into two distinct groups for disentangling content LoRA $\Delta W_{c}$ and style LoRA $\Delta W_{s}$.
B-LoRA identifies that the first block of Upblock0 mainly contributes for subject learning, while the second block of Upblock1 is more specialized for style learning. This observation holds when the input images are relatively simple or lack intricate details. However, when the input images have fine-grained details, only one block is insufficient to capture all the necessary information. To enhance the content learning, we utilize additional block as well. Moreover, we leverage all the blocks in the second group of the U-net upblocks for style learning in order to maintain a balance of the learning capacities between the content and style.
Reproduction and experimental settings. While reproducing the results of the competing methods, we use the exact hyperparameters reported in their respective papers (or their official implementations), including learning rate, training steps, and other experimental settings. For our approach, most experimental settings — such as the learning rate, batch size, and sampling frequency — are consistent with those described in the main paper. However, the number of training steps required for our method is $600$ for most input images, which is lesser as compared to other approaches. Depending on the complexity of the input however, our method may require higher number training steps (in the range of $800$ steps) if input image contains fine-grained details.
Training prompts. As mentioned in the main paper, we use three separate prompts during training: a combined prompt, and two separate prompts (for subject and style each). Our combined prompt is of the form “a <$c$> in <$s$> style”, while “A <$c$> ” and “An image in <$s$> style” are used as subject-specific and style-specific prompts respectively. Here, <$c$> and <$s$> represents the subject and style descriptors specific to the input image. We follow the descriptor strategies of existing approaches such as ZipLoRA [32] and StyleDrop [33] that advocates using a unique token (such as “sks”) to describe the subject, and use detailed artistic description for the style. For example, the prompt we use for the first input image in Fig. 12 is “A sks sun hat in flat cartoon illustration style”. As discussed in DreamBooth [29], using an unique token helps the model to bind the new subject to a unique embedding vector in space since a subject is usually spatially localized in a particular region of the image. On the other hand, style is typically spread throughout the image, thus using detailed text description increases the flexibility during style training as reported by StyleDrop [33]. To ensure a fair comparison between our method and others, such as B-LoRA, we use the same prompt formations for all the competing methods.
Figure 10: Ablation Examples. We show the impact of adding prompt-wise separation, column-wise separation, and block-wise separation sequentially. Each row illustrates the outputs of baseline methods and our proposed approaches, highlighting their difference in subject-style disentanglement, style fidelity, and recontextualization across different examples.
B Compute Requirements
UnZipLoRA achieves strong computational efficiency through a joint training strategy that optimizes resource utilization. By training both subject and style LoRAs concurrently in a single run, UnZipLoRA significantly reduces the overall training time. Specifically, UnZipLoRA requires only $1260$ seconds to train both LoRAs on a single NVIDIA A40 GPU.
In contrast, most existing methods necessitate separate training processes for content and style, effectively doubling the time requirements. For instance, DreamBooth-LoRA requires $1860$ seconds per LoRA, resulting in a total training time of $3720$ seconds. While B-LoRA demonstrates faster individual LoRA training at $600$ seconds per LoRA ($1200$ seconds total), UnZipLoRA remains highly competitive. Notably, methods like Inspiration Tree incur significantly higher computational costs, requiring $7680$ seconds in total: $3840$s to select a good random seed for training and another $3840$s to train the model.
Beyond time efficiency, UnZipLoRA minimizes the number of parameters updated during training. Through its block and column separation strategies, UnZipLoRA updates only up to $30\%$ of parameters in the downsampling block and bottleneck, and approximately $50\%$ in the upsampling block for each LoRA. This focused optimization reduces the trainable parameters by nearly $40\%$ compared to training two full LoRAs independently, further contributing to its efficiency. Owing to such efficient parameter utilization, UnZipLoRA exhibits faster convergence, requiring only $600$ steps of training as opposed to $800$ to $1000$ steps for most other methods including DreamBooth-LoRA and B-LoRA.
C Additional Results
C.1 Ablation Examples
We provide additional examples depicting the effects of adding each component of our method, as shown in Fig. 10. Each row demonstrates the recontextualization of subject, style and subject-style across different configurations and baseline methods.
Prompt separation. Distinct prompts and LoRA weights for subject and styles guide their respective LoRAs, ensuring effective and disentangled subject-style decomposition.
Column separation. Dynamic column masks selectively activate relevant columns during training, preserving learning capacity with fewer columns and preventing interference. Enhanced orthogonality between LoRAs improves flexibility in recombination.
Block separation. Style-sensitive blocks effectively capture essential stylistic features, while subject-sensitive blocks focus on fine details.
Figure 11: The interface used for user studies. We depict the interface we used for method comparison user study shown in the main paper.
Figure 12: Comparison of Subject-style Recontextualization. We present a comparison of subject-style recontextualization between our method and B-LoRA across diverse prompts and input images. The results highlight our method’s superior ability to flexibly adapt subjects and styles to various contexts while accurately reproducing both subject and style features.
We conducted user studies in the main paper to compare our method with the competing approaches. Beyond the configurations, results, and analyses presented in the main paper, we include the interface used for the main user study and the ablation user study in Fig. 11.
C.2 Subject-Style Recontextualization Comparison
We present additional examples of subject-style recontextualization using diverse prompts and input images comparing our method with B-LoRA in Fig. 12. The results further demonstrate our method’s superiority in preventing overfitting, reproducing accurate subject and style representations, and enabling flexible recontextualization.
Prevention of overfitting. Our method mitigates overfitting by disentangling subject and style representations, ensuring diverse and robust outputs even with challenging prompts
Accurate subject and style reproduction. We achieve precise reproduction of the input’s subject and style elements while avoiding blending artifacts.
Flexible recontextualization. Our method enables diverse and logical recontextualization, handling both straightforward prompts like ”on a table” and complex, creative prompts that require a nuanced extraction of subject and style.
Figure 13: Qualitative Comparisons. Additional groups of compare subject and style disentanglement from ours method against DreamBooth-LoRA, Inspiration Tree, and B-LoRA. The result again demonstrates our superior ability to preserve the intended features compared to other methods.
C.3 Additional Qualitative Comparisons
To complement the findings in the main paper, we provide additional qualitative comparisons in Fig. 13 across more diverse prompts and input images. These examples further demonstrate the superior performance of our method in subject-style disentanglement, detail preservation, and overfitting prevention compared to DreamBooth-LoRA, Inspiration Tree, and B-LoRA.
Upon closer inspection of the examples, our observations are consistent with the results presented in the main paper. Each of the compared methods demonstrates limitations that prevent them from achieving the level of disentanglement and flexibility required for our task.
DreamBooth-LoRA. DreamBooth-LoRA struggles to disentangle subject features from style, even though it captures some stylistic features effectively. However, its results often suffer from overfitting to the input image, limiting its ability to recontextualize the style in diverse settings. Our methods successfully captures the style with high fidelity, enabling flexible style recontextualization without overfitting to the input.
Inspiration Tree. While Inspiration Tree effectively prevents the overlap of subject and style concepts and consistently destylizes the input, it struggles to distinguish detailed features of both subject and style. This limitation results in outputs that lack the intricate details of the input subject or style. By incorporating separation strategies, our method intelligently learns and distinguishes these features, leading to more detailed and accurate outputs.
B-LoRA. As what we discussed in Experiments section of main paper, and in Sec. C.2, B-LoRA fails to generate consistent results and suffers from overfitting to input images. While it can always accurately learn style features, it struggles to reproduce subject details reliably. For instance, in the guitar and sunflower examples, B-LoRA fails to consistently retain the original input’s color.
Moreover, it often mixes subject and style features, resulting in generations that incorporate unintended elements, such as the background color of the microscope or the sunflower’s color being treated as part of the style. In contrast, our method addresses these issues with carefully designed separation techniques, ensuring consistent, disentangled outputs that faithfully represent both subject and style.
D Image Attributions
In our experiments, we use several stylized images as inputs images. We curate these input images from three sources: (i) free-to-use online repositories that provide artistic images; (ii) open-sourced repositories of previous works such as StyleDrop [33] and RB-Modulation [28]; and (iii) synthetically generated images using freely available text-to-image models such as Flux. For the human-created artistic images, we provide image attributions below for each image that we used in our experiments.
D.1 Image attributions for the stylized images used as inputs
All the remaining input images are generated using Flux [^2] https://huggingface.co/black-forest-labs/FLUX.1-dev text-to-image diffusion model using the prompts provided by RB-modulation codebase on their github page at this URL: Text-prompts to generate stylized images